OpenStudy (anonymous):

PLEASEE HELP MEEE!:/ What is the sum of the arithmetic sequence 149,135,121,..., if there are 28 terms? a. -1,204 b.-1,176 c. -1,148 d. -1,120

OpenStudy (anonymous):

keep subtracting 14 i think

OpenStudy (anonymous):

i did but then it become negative & gets lower:/ get me?

OpenStudy (anonymous):

the ancer is D -1220

OpenStudy (anonymous):

answer

OpenStudy (marigirl):

yes D i the answer

OpenStudy (marigirl):

S = (n/2) × (2a + (n-1)d) This is the formula s=(28/2)*((2*149)+(28-1)*-14)=-1120

OpenStudy (marigirl):

sorry n stands for the number of terms in the sequence, a- first term d- difference

OpenStudy (anonymous):

thank you so much!